OpenStudy (anonymous):

find the equation of tangent to graph y=e^x at point (0,1)

OpenStudy (anonymous):

frist find slope of tangent (m=dy/dx )at point (0,1) since dy/dx=e^x so dy/dx at (0,1) is = e^0=1 so m=1 now use equation os straight line y=mx+c, put value m=1 eqution becomes y=x+c now put x=0 and y=1 in above eqution you will get c which is =1 so finally you will get equation of tangent which is y=x+1